Blast it! said David Wilson, president of Teledex Company. Weve just lost the bid on the Koopers job by $4,000. It seemsReferences b. What would the bid price have been if departmental predetermined overhead rates had been used to apply overhead
Answer #1

Predetermined Overhead Rate = Estimated overheads/estimated direct labor cost

= 907200/648000 = 140.00%

Applied on Koopers Job

             = 14300*140% = $20,020

Fabricating

Machining

Assembly

Manufacturing Overhead

            378,000

                      432,000

97,200

Direct Labor

216,000

108,000

324,000

Predetermined Rate

175%

400%

30%

b. Applied to Koopers Job

            = 6000*175% + 500*400% + 7800*30% = $14,840

Plantwide Rate

Departmental Rates

Direct Material

                 7,800

                          7,800

Direct Labor

14,300

14,300

Overheads

20,020

14,840

Total cost

42,120

36,940

Bid Price

              63,180

                        55,410